A sample of nitrogen gas exerts a pressure of 9.80 atm at 32 °C. What would its temperature be (in °C) when its pressure is increased to 11.2 atm?
1
Expert's answer
2013-02-28T11:14:20-0500
For answering this question you need to use the next law:
PV = nRT
V, R, n are constants
P is pressure
T is temperature
From the first law:
P1/T1 = P2/T2
T2 = P2/(P1/T1)
So the answer is T2 = 36.6